OpenStudy (anonymous):

cot^2 x-csc^2 x=-1 true or false

OpenStudy (whpalmer4):

\[\cot ^2(x)-\csc ^2(x) = \frac{\cos^2(x)}{\sin^2(x)} - \frac{1}{\sin^2(x)} = \]

OpenStudy (whpalmer4):

If \[\sin^2(x) + \cos^2(x) = 1\]what does \(\cos^2(x)-1=\)

OpenStudy (anonymous):

i did that already but still thank you :)